Shift report 20030224 00:00-08:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Mon Feb 24 2003 - 07:52:30 EST

  • Next message: Apache: "Shift report 20030224 08:00-16:00"
    ------------------------------------------------------------
    Date: Mon Feb 24 07:52:30 2003, Shift: 00:00-08:00
    Supervisor: BHS, with: 
    Log book#14 pages 34 - 34
    
    Shift Summary:
    ------------------------------------------------------------
    1: RHIC has tried three stores tonght, two of them failing after a short time (minutes...) and the third one just getting under ways now. Common for all of them are huge backgrounds in our IR, peaking at a TOFW rate of 1.5M for the  second store...
    
    2: Only pedestal/DVM runs tonight.
    
    3: No problems, though. TPM1 and TPM2 are still OFF due to the high BG.
    
    Just now, MCR tried to "optimize beam conditions" and our TOFW rate went from 200k to 400k. I won't start a run under these conditions, but will rather leave it to the next shifter to request more steering if possible...
    
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Mon Feb 24 2003 - 07:53:45 EST